# Location–Scale Calibration for Generalized Posterior

### v1

@@ Abstract
General Bayesian updating replaces the likelihood with a loss scaled by a learning rate, but posterior uncertainty can depend sharply on that scale. We propose a simple post-processing that aligns generalized posterior draws with their asymptotic target, yielding uncertainty quantification that is invariant to the learning rate. We prove total-variation convergence for generalized posteriors with an effective sample size, allowing sample-size-dependent priors, non-i.i.d. observations, and convex penalties under model misspecification. Within this framework, we justify and extend the open-faced sandwich adjustment (Shaby,, 2014), provide general theoretical guarantees for its use within generalized Bayes, and extend it from covariance rescaling to a location–scale calibration whose draws converge in total variation to the target for any learning rate. In our empirical illustration, calibrated draws maintain stable coverage, interval width, and bias over orders of magnitude in the learning rate and closely track frequentist benchmarks, whereas uncalibrated posteriors vary markedly.
@@ 1 Introduction
Bayesian inference provides a coherent probabilistic framework that combines prior information with likelihood-based learning and delivers uncertainty quantification. However, when the assumed likelihood is misspecified, or when inference is based on a non-likelihood objective such as quasi- or composite likelihoods, estimating equations, or other loss-based objectives, posterior uncertainty can be miscalibrated (Kleijn and van der Vaart,, 2012; Syring and Martin,, 2019; Miller,, 2021). General Bayesian inference replaces the likelihood with a loss-based construction (Bissiri et al.,, 2016), thereby avoiding the need to model the entire data distribution explicitly. In this formulation, a single scaling parameter $\eta\in\mathbb{R}_{>0}$, often called the learning rate or temperature, multiplies the loss and directly controls posterior dispersion. When $\eta=1$ and the loss function is taken as the negative log-likelihood, the generalized posterior coincides with the usual likelihood-based Bayes posterior. With a fixed prior, decreasing $\eta$ attenuates the contribution of the data so that, in the limit, the update reverts to the prior, whereas increasing $\eta$ amplifies the data contribution and, under standard regularity conditions, the posterior concentrates around minimizers of the population loss. Thus, the learning rate governs the trade-off between prior information and loss-based evidence.
@@
Therefore, the choice of learning rate has been widely discussed. Bootstrap-based calibration methods choose $\eta$ to achieve frequentist targets (Lyddon et al.,, 2019; Syring and Martin,, 2019); in a related approach, Matsubara et al., (2024) compute the loss minimizer on bootstrap resamples, obtain a closed-form estimate of the learning rate from the bootstrap spread, and then run Markov chain Monte Carlo. SafeBayes provides an alternative data-driven choice that adapts the learning rate for robustness (Grünwald and van Ommen,, 2017). Information-matching rules select the learning rate so that the generalized posterior with learning rate $\eta$ aligns, under criteria of divergence or information, with the generalized posterior obtained by setting $\eta=1$ for the same loss (Holmes and Walker,, 2017). Along a related calibration-to-Bayes line, Altamirano et al., (2023) choose the learning rate by minimizing the Kullback–Leibler divergence between the generalized posterior with learning rate $\eta$ and the generalized posterior obtained by setting $\eta=1$ for the same loss, both computed on an initial data window, and then keep the resulting learning rate fixed for the full analysis. These approaches can be computationally intensive and typically require a frequentist point estimator, so the resulting learning rate reflects the plug-in distribution of that estimator rather than the full posterior law. Moreover, several methods explicitly or implicitly calibrate to the generalized posterior obtained by setting $\eta=1$ for the chosen loss; this is natural when the loss is negative log-likelihood under correct specification, but can be misleading for more general loss-based objectives or under model misspecification. Although McLatchie et al., (2025) suggest that, for prediction, the choice of learning rate matters little in moderate to large samples, this does not resolve the problem of the sensitivity of uncertainty quantification to the learning rate.
@@
In this paper, we revisit the open-faced sandwich adjustment (Shaby,, 2014) and place it in a general asymptotic framework for generalized Bayes based on loss functions with an effective sample size, allowing sample-size-dependent priors, possibly non-i.i.d. observations, and convex penalties under model misspecification. First, we show that, under sample-size-dependent priors, the generalized posterior admits a normal limit in total variation and we establish a prior–penalty correspondence that identifies the target curvature and variability. Second, building on this justification, we improve the open-faced sandwich adjustment from covariance rescaling to a location–scale calibration that is implementable via plug-in sandwich estimators, requiring only posterior draws and empirical moments (no bootstrap or learning-rate tuning). Third, we prove that the calibrated draws converge to the target distribution for any learning rate, so that the asymptotic uncertainty quantification is invariant to the choice of learning rate. Our empirical illustration for a random-intercept mixed model with Huber loss and Gaussian prior with sample-size-dependent scale shows that the calibrated intervals track frequentist benchmarks across several orders of magnitude in the learning rate, whereas uncalibrated generalized Bayes posteriors exhibit substantial sensitivity.
@@ 2 Problem setup
For a learning rate $\eta>0$, we adopt the general Bayesian updating framework of Bissiri et al., (2016). Define the composite loss $L_{n}(\boldsymbol{\theta}):=M_{n}(\boldsymbol{\theta})+\lambda_{n}s_{n}\rho(\boldsymbol{\theta})$, where $\{\lambda_{n}\}_{n\geq 1}$ is a deterministic non-negative sequence with $\lambda_{n}\to\lambda\in[0,\infty)$ and $\rho:\mathbb{R}^{p}\to\mathbb{R}$ is a convex penalty. We introduce a baseline prior $\tilde{\pi}_{n}(\boldsymbol{\theta})\propto\exp\{r_{n}(\boldsymbol{\theta})\}$ that may depend on $n$ but is independent of the data. The generalized posterior is then given by
@@
This loss-based representation makes explicit that the learning rate $\eta$ scales the entire loss, in line with Bissiri et al., (2016). Throughout the main text we assume that $\rho$ is $C^{2}$ in a neighborhood of the target point introduced below (see Assumption 2.1); the convex non-smooth case is treated in Section D.
@@
The matrix $\boldsymbol{V}_{\mathrm{target}}^{\star}$ is the usual sandwich covariance for penalized $M$-estimators; see, e.g., van der Vaart, (1998); Kosorok, (2008).
@@ 3 Main results
We denote by $q_{n}^{\tilde{\boldsymbol{\theta}}_{n}}$ the density of $\sqrt{s_{n}}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$ under $\Pi_{n}^{\eta}$, and by $\mathcal{N}(\cdot\mid\boldsymbol{\mu},\boldsymbol{\Sigma})$ the density of the $p$-variate normal distribution with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$. Using local asymptotic normality arguments in the spirit of Miller, (2021, Theorem 4), adapted to our loss-based criterion with effective scale $s_{n}$, to possibly $n$-dependent priors and to non-i.i.d. data, and evaluated at admissible centers, we obtain the following total-variation limit.
@@ 5.2 Experiment
We illustrate the finite-sample behavior of the plug-in location–scale calibration in the random-intercept linear mixed model with Huber loss and Gaussian prior with $n$-dependent scale described in Section 5.1. We fix $G=100$, $n_{i}=5$ (so $n=500$), and set $p=1$. We generate covariates from $\mathcal{N}(0,1)$ and contaminate Gaussian errors to induce model misspecification. Since no closed form is available for $\beta^{\lambda}$ in this example, we approximate it numerically by computing the penalized estimating equation estimator on large simulated datasets with $G=5,000$ and averaging over $1,000$ replications; see, e.g., Oh and Patton, (2013). For each learning rate $\eta$ on a logarithmic grid over $[0.01,100]$ we compare three procedures: (i) a frequentist Huber $M$-estimator with ridge penalty and sandwich-based Wald intervals; (ii) the generalized Bayes posterior based on $M_{n}$ and the corresponding $n$-dependent Gaussian prior; and (iii) the location–scale calibration applied to the same posterior draws, using the MAP estimate as an admissible center. For each $\eta$ we use $200$ Monte Carlo replications and record, for all three methods, the empirical coverage, the mean interval width, the mean bias, and the standard deviation of that bias. Further details of the data-generating mechanism, the numerical approximation of $\beta^{\lambda}$, and the Markov chain Monte Carlo settings are provided in Section E. The Python code for reproducing the experiments is available at https://github.com/shutech2001/ls-calib-gp.
@@ Appendix B Proofs of the main results
For the prior remainder $r_{n}$, Assumption 2.3 states that $r_{n}$ is locally Lipschitz on $\mathcal{U}_{n}$ with Lipschitz constant $L_{n}=o_{p}(\sqrt{s_{n}})$. Lebourg’s mean value theorem for locally Lipschitz functions (see, e.g., Clarke, (1990, Thm. 2.3.7)) applied along the segment joining $\check{\boldsymbol{\theta}}_{n}$ and $\check{\boldsymbol{\theta}}_{n}+\boldsymbol{u}/\sqrt{s_{n}}$ gives, for each fixed $M<\infty$ and all $\|\boldsymbol{u}\|\leq M$,

### v2

@@ Abstract
General Bayesian updating replaces the likelihood with a loss scaled by a learning rate, but posterior uncertainty can depend sharply on that scale. We propose a simple post-processing that aligns generalized posterior draws with their asymptotic target, yielding uncertainty quantification that is invariant to the learning rate. We prove total-variation convergence for generalized posteriors with an effective sample size, allowing sample-size-dependent priors, non-i.i.d. observations, and convex penalties under model misspecification. Within this framework, we justify and extend the open-faced sandwich adjustment (shaby2014open-faced), provide general theoretical guarantees for its use within generalized Bayes, and extend it from covariance rescaling to a location–scale calibration whose draws converge in total variation to the target for any learning rate. In our empirical illustration, calibrated draws maintain stable coverage, interval width, and bias over orders of magnitude in the learning rate and closely track frequentist benchmarks, whereas uncalibrated posteriors vary markedly.
@@ 1 Introduction
Bayesian inference provides a coherent probabilistic framework that combines prior information with likelihood-based learning and delivers uncertainty quantification. However, when the assumed likelihood is misspecified, or when inference is based on a non-likelihood objective such as quasi- or composite likelihoods, estimating equations, or other loss-based objectives, posterior uncertainty can be miscalibrated (kleijn2012bernstein; syring2019calibrating; miller2021asymptotic). General Bayesian inference replaces the likelihood with a loss-based construction (bissiri2016general), thereby avoiding the need to model the entire data distribution explicitly. In this formulation, a single scaling parameter $\eta\in\mathbb{R}_{>0}$, often called the learning rate or temperature, multiplies the loss and directly controls posterior dispersion. When $\eta=1$ and the loss function is taken as the negative log-likelihood, the generalized posterior coincides with the usual likelihood-based Bayes posterior. With a fixed prior, decreasing $\eta$ attenuates the contribution of the data so that, in the limit, the update reverts to the prior, whereas increasing $\eta$ amplifies the data contribution and, under standard regularity conditions, the posterior concentrates around minimizers of the population loss. Thus, the learning rate governs the trade-off between prior information and loss-based evidence.
@@
Therefore, the choice of learning rate has been widely discussed. Bootstrap-based calibration methods choose $\eta$ to achieve frequentist targets (lyddon2019general; syring2019calibrating); in a related approach, matsubara2024generalized compute the loss minimizer on bootstrap resamples, obtain a closed-form estimate of the learning rate from the bootstrap spread, and then run Markov chain Monte Carlo. SafeBayes provides an alternative data-driven choice that adapts the learning rate for robustness (grunwald2017inconsistency). Information-matching rules select the learning rate so that the generalized posterior with learning rate $\eta$ aligns, under criteria of divergence or information, with the generalized posterior obtained by setting $\eta=1$ for the same loss (holmes2017assigning). Along a related calibration-to-Bayes line, altamirano2023robust choose the learning rate by minimizing the Kullback–Leibler divergence between the generalized posterior with learning rate $\eta$ and the generalized posterior obtained by setting $\eta=1$ for the same loss, both computed on an initial data window, and then keep the resulting learning rate fixed for the full analysis. These approaches can be computationally intensive and typically require a frequentist point estimator, so the resulting learning rate reflects the plug-in distribution of that estimator rather than the full posterior law. Moreover, several methods explicitly or implicitly calibrate to the generalized posterior obtained by setting $\eta=1$ for the chosen loss; this is natural when the loss is negative log-likelihood under correct specification, but can be misleading for more general loss-based objectives or under model misspecification. Although mclatchie2025predictive suggest that, for prediction, the choice of learning rate matters little in moderate to large samples, this does not resolve the problem of the sensitivity of uncertainty quantification to the learning rate.
@@
In this paper, we revisit the open-faced sandwich adjustment (shaby2014open-faced) and place it in a general asymptotic framework for generalized Bayes based on loss functions with an effective sample size, allowing sample-size-dependent priors, possibly non-i.i.d. observations, and convex penalties under model misspecification. First, we show that, under sample-size-dependent priors, the generalized posterior admits a normal limit in total variation and we establish a prior–penalty correspondence that identifies the target curvature and variability. Second, building on this justification, we improve the open-faced sandwich adjustment from covariance rescaling to a location–scale calibration that is implementable via plug-in sandwich estimators, requiring only posterior draws and empirical moments (no bootstrap or learning-rate tuning). Third, we prove that the calibrated draws converge to the target distribution for any learning rate, so that the asymptotic uncertainty quantification is invariant to the choice of learning rate. Our empirical illustration for a random-intercept mixed model with Huber loss and Gaussian prior with sample-size-dependent scale shows that the calibrated intervals track frequentist benchmarks across several orders of magnitude in the learning rate, whereas uncalibrated generalized Bayes posteriors exhibit substantial sensitivity.
@@ 2 Problem setup
For a learning rate $\eta>0$, we adopt the general Bayesian updating framework of bissiri2016general. Define the composite loss $L_{n}(\boldsymbol{\theta}):=M_{n}(\boldsymbol{\theta})+\lambda_{n}s_{n}\rho(\boldsymbol{\theta})$, where $\{\lambda_{n}\}_{n\geq 1}$ is a deterministic non-negative sequence with $\lambda_{n}\to\lambda\in[0,\infty)$ and $\rho:\mathbb{R}^{p}\to\mathbb{R}$ is a convex penalty. We introduce a baseline prior $\tilde{\pi}_{n}(\boldsymbol{\theta})\propto\exp\{r_{n}(\boldsymbol{\theta})\}$ that may depend on $n$ but is independent of the data. The generalized posterior is then given by
@@
This loss-based representation makes explicit that the learning rate $\eta$ scales the entire loss, in line with bissiri2016general. Throughout the main text we assume that $\rho$ is $C^{2}$ in a neighborhood of the target point introduced below (see Assumption 2.1); the convex non-smooth case is treated in Section D.
@@
The matrix $\boldsymbol{V}_{\mathrm{target}}^{\star}$ is the usual sandwich covariance for penalized $M$-estimators; see, e.g., vandervaart1998asymptotic; kosorok2008introduction.
@@ 3 Main results
We denote by $q_{n}^{\tilde{\boldsymbol{\theta}}_{n}}$ the density of $\sqrt{s_{n}}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$ under $\Pi_{n}^{\eta}$, and by $\mathcal{N}(\cdot\mid\boldsymbol{\mu},\boldsymbol{\Sigma})$ the density of the $p$-variate normal distribution with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$. Using local asymptotic normality arguments in the spirit of miller2021asymptotic, adapted to our loss-based criterion with effective scale $s_{n}$, to possibly $n$-dependent priors and to non-i.i.d. data, and evaluated at admissible centers, we obtain the following total-variation limit.
@@ 5.2 Experiment
We illustrate the finite-sample behavior of the plug-in location–scale calibration in the random-intercept linear mixed model with Huber loss and Gaussian prior with $n$-dependent scale described in Section 5.1. We fix $G=100$, $n_{i}=5$ (so $n=500$), and set $p=1$. We generate covariates from $\mathcal{N}(0,1)$ and contaminate Gaussian errors to induce model misspecification. Since no closed form is available for $\beta^{\lambda}$ in this example, we approximate it numerically by computing the penalized estimating equation estimator on large simulated datasets with $G=5,000$ and averaging over $1,000$ replications; see, e.g., oh2013simulated. For each learning rate $\eta$ on a logarithmic grid over $[0.01,100]$ we compare three procedures: (i) a frequentist Huber $M$-estimator with ridge penalty and sandwich-based Wald intervals; (ii) the generalized Bayes posterior based on $M_{n}$ and the corresponding $n$-dependent Gaussian prior; and (iii) the location–scale calibration applied to the same posterior draws, using the MAP estimate as an admissible center. For each $\eta$ we use $200$ Monte Carlo replications and record, for all three methods, the empirical coverage, the mean interval width, the mean bias, and the standard deviation of that bias. Further details of the data-generating mechanism, the numerical approximation of $\beta^{\lambda}$, and the Markov chain Monte Carlo settings are provided in Section E. The Python code for reproducing the experiments is available at https://github.com/shutech2001/ls-calib-gp.
@@ Appendix B Proofs of the main results
For the prior remainder $r_{n}$, Assumption 2.3 states that $r_{n}$ is locally Lipschitz on $\mathcal{U}_{n}$ with Lipschitz constant $L_{n}=o_{p}(\sqrt{s_{n}})$. Lebourg’s mean value theorem for locally Lipschitz functions (see, e.g., clarke1990optimization) applied along the segment joining $\check{\boldsymbol{\theta}}_{n}$ and $\check{\boldsymbol{\theta}}_{n}+\boldsymbol{u}/\sqrt{s_{n}}$ gives, for each fixed $M<\infty$ and all $\|\boldsymbol{u}\|\leq M$,